OpenStudy (anonymous):

MAX h(x) = f(x) - g(x) f(x) = 29x – x^2 g(x) = x^2 + x + 10

OpenStudy (john_es):

First find h(x), \[h(x)=29x-x^2-(x^2+x+10)\]Simplify. Then find the derivative, h'(x) and do, \[h'(x)=0\]Solve for x, and study the sign of h'(x).

OpenStudy (anonymous):

Iam getting h(x)=-2x^2+28x-10 h´(x)=-4x+28 h'(x) = how do you get the (-6x)

OpenStudy (john_es):

Because I confused a 2 with a 3 in the power. You are correct!! \[h(x)=−2x^2+28x−10\\ h'(x)=-4x+28 \\-4x+28=0\Rightarrow x=-7\]